# An Optimal, Universal and Agnostic Decoding Method for Message Reconstruction, Bio and Technosignature Detection

### v3

@@ 3 On the information content of the Arecibo message
By translating (encoding) a colour image into binary black and white pixels, this method picks an image decoding that is similar enough to the original image (at least similar enough visually). Fig. LABEL:mandril shows the original image on the left, the results of the method on the right, and the resulting image selected from the spike in BDM on the bottom. A mirror-like image is selected as the first candidate, followed by the correct one. This kind of spiking also suggests the original partition can be inferred via smaller, non-random spikes at multiples or divisors of the native partition (500). In this case, small spikes at half the original partition (250) provide clues to the original embedding in 2D.
4 Discussion
@@
This work advances a practical and theoretical framework that relates information, entropy, complexity, and semantics that can be extended and hence put to multiple uses in signal deconvolution, bio- and technosignature detection, cryptography, and coding theory. Our methods show how the receiver can decode the (multidimensional) space into which the original message was sent via zero-knowledge one-way communication channels.
@@ Acknowledgements
Author Contributions: HZ conceived the theory, methods and experiments; HZ performed most of the experiments, with support from AA. HZ and FA developed the theoretical framework.
Appendix A Figures
Figure 1: Algorithmic perturbation analysis on short sentences of only 400 characters from Darwin’s “Origin of Species” are converted from ASCII into their corresponding binary values. Changing a single character with probability $1/n$ for growing $n$ from $n=1$ (original test, axis $x=0$) to $n=4$ (a quarter of the text is mapped to a random character from the same vocabulary of the original text, axis $x=100$ letters changed or perturbed) causes the resulting size of the compressed sequence index to increase, with BDM the most sensitive and Shannon entropy unable to capture any difference. This is because signals carrying meaning are far removed from randomness, and random perturbations make the text more random [35] even when the methods know nothing about words, grammar or anything linguistic.
Figure 2: Top: Sound waves of the words spoken from Apollo 13, “Houston, we’ve had a problem.” transmitted on April 13, 1970 at a sample rate frequency of 11025 Hz on a single channel. Bottom left: Histograms of the change in compression lengths of scrambled versions of the same message from the length of the original compressed message. Bottom right: Small perturbations of the original message by scrambling only the beginning and end of each word shows a smooth transition from low to high randomness. The lowest complexity signal indicates the correct (original) signal. The file was processed in FLAC format (Free Lossless Audio Codec) from a lossless file, with no audio data discarded.
Refer to caption           Refer to caption
Figure 4: Top left: Most possible partitions result in random-appearing configurations with high corresponding complexity, indicating measurable randomness. Bottom: Some partitions will approximate the originally encoded meaning (third from the right). Other configurations result in images with higher complexity values. This sequence of images shows the images in the approximate vicinity of the correct bidimensional configuration (i.e., partition) and illustrates fast convergence to low complexity. Top right: By using different information indexes across different configurations, a downward-pointing spike will indicate message (image) configurations that correspond to low-complexity image(s). This allows a prior-knowledge-agnostic and objective method to infer a message’s original encoding. Of the various measures, BDM, combining classical information (entropy) for long ranges and a measure motivated by algorithmic probability for short ranges, is the most sensitive and accurate in this regard. Traditional compression and entropy also indicate the right configuration amongst the top spiking candidates. The ratio of noise-to-signal was amplified in favour of the hidden structure by multiplying the original image size by 6 for both length and height.
Figure 6: Applying 3 different quantitative methods, it is shown (left) that they are highly insensitive to signal and highly sensitive to noise (16.5% of pixels randomly flipped). However, by growing the original image (right) by a factor of 6 on each dimension (i.e. 1 pixel becomes $6\times 6$), the methods are less sensitive to noise and more sensitive to signal, with BDM significantly outperforming Compression, and Shannon Entropy showing sensitivity at up to 60% pixels flipped (hence about 30% of the original image) versus Compression and Shannon Entropy, that are about 50% sensitive. Downward spikes (right) are shown at $23\times 6=141$.
196                                   82                                     215
276                                 360                                 220
Figure 7: Six 2D images (labelled 1 through 6 going from left to right, top to bottom) of very different nature, including a demonstration of linear-transformation invariance conforming to the underlying theory (in this case, rotation of a mathematical formula). Size invariance has actually shown amplification of signal-to-noise difference. Under each image is its correct numerical (first) dimension. The values of the three complexity indexes over possible partitions are shown below the original numerical dimension for ease of comparison. Downward spikes indicate candidates for possible original partitions. In all cases, the correct dimension value is among the top three candidates, with BDM outperforming in 4 out of 6 cases at indicating the top candidate.
Figure 8: Sequence of lowest complexity (BDM value on top in bits) partitions of a black and white image of a space shuttle, with the top right partition as the correct reconstruction and other candidates cases in which an alignment occurs at some multiple of the correct dimension.

### v4

@@ 2 On the information content of the Arecibo message
In translating (encoding) a colour image into binary black and white pixels, this method picks an image decoding that is similar enough to the original image (at least similar enough visually). Fig. LABEL:mandril shows the original image on the left, the results of the method on the right, and the resulting image selected from the spike in BDM on the bottom. A mirror-like image is selected as the first candidate, followed by the correct one. This kind of spiking also suggests that the original partition can be inferred via smaller, non-random spikes at multiples or divisors of the native partition (500). In this case, small spikes at half the original partition (250) provide clues to the original encoding in 2D.
3 Assessing the method’s performance over several different classes of images
So far, we have illustrated the applicability of the method to a few examples of image data. In order to investigate its generalisability, we considered 102 images taken from the Caltech 101 dataset [38]. Described briefly, this dataset contains pictures of objects belonging to 101 categories and an extra category named background, and the size of each image is roughly 300 x 200 pixels.
In particular, we will test the reshaping part of the method, to check if it is capable of shortlisting the correct dimensions of each image for such a diverse set. As detailed in [26], we consider the upscaling of the arrays (changing a single bit to a square of size $6\times 6$—therefore a resulting $36x$ total change in the image size $x$). This is to make sure z-lib analysis will not be impaired by small-size samples.
Since there are several rearrangements possible, it was observed for every case studied so far that only the downward peak values are of interest. In particular, we filter only the downward peaks of BDM and its immediate left and right neighbours, selecting them if the normalised BDM value change between them and the downward peak is at most $1/10$. This is done to put to the test the hypothesis that for every image, the correct dimension will be either a downward peak or its immediate neighbours.
After running the method on the 102 images, a total of 29511 downward BDM peaks and eligible neighbours was observed. In Fig. 9 the values of (BDM,Entropy,z-lib) are presented for each candidate peak, which are marked as either blue or red. The correct dimensions are the ones whose metrics are marked as red squares, while the blue circles are candidate peaks which were not correct. A very low alpha was chosen for the blue circles in order to get a clear view of the red squares.
Figure 9: 3D plot of the values the of three metrics for all the downward peaks and selected neighbours. Red squares, which indicate the correct dimensions of the input message, are clearly clustered in the low-normalised BDM, low-normalised compressibility (high z-lib values) and high-normalised entropy region, revealing that these three metrics, when used together, can select good candidates for such geometrical features of the input message.
Actually, from Fig. 9 it is possible to see that $80\%$ of the correct dimensions are the ones where the lowest BDM value is observed. Some border effects that may show up when applying BDM (which will be studied in detail in subsequent papers), could perhaps account for the fact that the other 20 % of cases did not show up as the lowest BDM configuration. If we consider a decision tree of depth 6 to classify the peaks as either correct (1) or incorrect (0), all the peaks can be successfully separated. This decision tree is presented in the Sup. Inf..
It is important to highlight that the studies herein presented are a first approach to the problem, such that the decision tree above could not be used to extrapolate to any other image. On the other hand, it is clear that the method proposed is able to cluster candidates with correct dimensions, which empirically verifies its applicability and overall correctness.
Future studies could benefit from checking if any other metrics (or the combination of the ones already used) would be better suited to the task. As indicated, sometimes the correct dimension may not be the exact tip of the downward peak, but its immediate right or left neighbour. This may result from slight changes in the image that are not properly reflected on BDM due to the large size of the image when compared to the $4\times 4$ grid used to partition it prior to BDM calculation. Also, it could be of interest to analyse which peaks to select based on their width and prominence.
The experiments performed also revealed that when dealing with silhouettes of images, the candidate peaks are easier to classify (mostly because the background is filled with 0s). When full real images come into play, then deeper decision trees are needed.
4 On the link between the universal distribution, AID, the method hereby proposed and Artificial General Intelligence (AGI)
The universal distribution, also known as Levin’s distribution or the Solomonoff-Levin distribution, plays a central role in algorithmic information theory (AIT) [39, 40, 41, 42] by describing the optimal/maximal probability of arbitrarily generating an encodable object from a randomly generated program running on any universal Turing machine. This distribution is portrayed as a way to capture the likelihood of patterns or data occurring naturally in our universe, reflecting the intrinsic complexity (or simplicity) of objects [33]. The simpler a program capable of generating a particular string, the higher its probability under the universal distribution, which is maximal for any arbitrary formal-theoretic probability measure of the space of computably constructible objects, and therefore demonstrating its depth and pervasiveness in mathematics and science.
The connection between the universal distribution and AGI lies in the way AGI systems need to process and learn from diverse, complex environments. AGI is expected to generalise learning across different domains, in much the same way that the universal distribution captures the probability of generating any pattern or object. By relying on algorithmic probability, which is one of underpinnings of the universal distribution, AGI systems can prioritise more likely solutions, especially when dealing with unknown or vast search spaces. The universal distribution aligns with AGI’s requirement to balance exploration and exploitation—i.e., finding efficient solutions (simple programs with high probabilities) while remaining open to novel possibilities (complex patterns with lower probabilities). This framework supports AGI’s ability to solve new, unseen problems with minimal data, mimicking human intuition, which tends to favour simpler, concise, non-redundant, and (equivalently, as demonstrated in AIT) more likely explanations based on prior experience.
In essence, the universal distribution provides a mathematical grounding for AGI’s learning mechanisms by offering a way to estimate the likelihood of different hypotheses or patterns, thus guiding the system toward efficient learning and decision-making. In this scenario, the method here proposed builds on AID and on the universal distribution, using such theoretical frameworks to explore how information and complexity evolve across both space and time. AID, an extension of AIT, moves beyond algorithmic complexity by emphasising how systems dynamically adapt and transform within structured or random contexts. These transformations were assessed by applying perturbations to both individual elements of an input message and to its overall geometric configuration. Our present paper also extends AIT by not relying on the choice of a fixed mathematical space or structure [26]. Such an approach is highly relevant to AGI, as AGI systems must demonstrate the ability to flexibly adapt to novel, unforeseen challenges without prior domain-specific knowledge.
By utilising tools like the Coding Theorem Method (CTM) and the Block Decomposition Method (BDM), we can more precisely evaluate the complexity of incoming data and gain deeper insights into how an original emitter processes, stores, and generates information. AGI requires similar competencies, such as generalisation, adaptability, and the ability to learn from minimal information. Thus, both AID and the here proposed method align naturally with the core challenges of AGI research, offering a framework for intelligent systems to navigate complexity efficiently.
Moreover, the relationship between algorithmic complexity and intelligence is crucial, as AGI systems must manage complex information effectively. Techniques like BDM, which break down and assess the complexity of structured and unstructured data, provide a quantitative approach to how intelligent systems should approach learning and problem-solving. This dynamic understanding of complexity aids in the design of AGI systems capable of evolving their strategies, ensuring that they can not only handle static tasks but also adjust flexibly, mirroring human adaptability and intelligence.
5 Conclusions
@@
This work advances a practical and theoretical framework that relates information, entropy, complexity, and semantics that can be extended and hence put to multiple uses in signal deconvolution, bio- and technosignature detection, cryptography, and coding theory. Our methods show how the receiver can decode the (multidimensional) space or structure into which the original message was sent via zero-knowledge one-way communication channels.
@@ Acknowledgements
Author Contributions: HZ conceived the theory, methods and experiments; HZ performed most of the experiments, with support from AA and LO. HZ and FA developed the theoretical framework.